OpenStudy (toxicsugar22):

How much heat energy is required to raise the temperature of 0.358kg of copper from 23.0 ∘C to 60.0 ∘C? The specific heat of copper is 0.0920 cal/(g⋅∘C) .

OpenStudy (anonymous):

this is physic hehe

OpenStudy (anonymous):

lol i will try to solve it

OpenStudy (anonymous):

is it 1.3105k cal?

OpenStudy (anonymous):

\[C = E/(m \Delta T)\] C = 0.092, m = 358g, delta T = 37 \[E =C(m \Delta T) = 0.092 * 358 * 37 = 1.2186 k \cal\] i think i have copied the wrong number in last post

OpenStudy (shrutipande9):

i guess that would be \[Q=mc \Delta t\]t m=0.358kg=358g c=0.0920 cal/g deg C \[\Delta t\]=60-23=37 deg C Q=358*0.0920*37=1218 cal

OpenStudy (anonymous):

yeah 1218, or 1.218k cal

OpenStudy (shrutipande9):

energy in cal or J

OpenStudy (shrutipande9):

its just that you need to convert the mass of copper in g because rest all the values are in CGS system...

OpenStudy (shrutipande9):

but is the ans correct?

OpenStudy (shrutipande9):

sure

OpenStudy (toxicsugar22):

If 125 cal of heat is applied to a 60.0-g piece of copper at 20.0∘C , what will the final temperature be? The specific heat of copper is 0.0920 cal/(g⋅∘C) .

OpenStudy (shrutipande9):

the same formula goes here \[Q=mc Deltat\] 125=60*0.0920*(x-20) 125=5.52*(x-20) 22.65=x-20 x=42.65

OpenStudy (shrutipande9):

correct?

OpenStudy (shrutipande9):

yes yes...temp in C

OpenStudy (shrutipande9):

ur welcum:)

OpenStudy (toxicsugar22):

hwy if you can help me with one more i would appreciat it

OpenStudy (toxicsugar22):

A cyclist rides at an average speed of 28mi/h .

OpenStudy (toxicsugar22):

If she wants to bike 181km , how long (in hours) must she ride?

OpenStudy (toxicsugar22):

hei can you help me

OpenStudy (toxicsugar22):

i got 4 hours is that right

OpenStudy (shrutipande9):

speed=distance/time use this formula...dat mi is miles right? i dnt know miles to km conversion. convert it and then just put the values in this formula.

OpenStudy (shrutipande9):

convert km to miles...dat would be easy

OpenStudy (toxicsugar22):

i got 4 hours is it it right

OpenStudy (shrutipande9):

28=112.47/t t=112.47/28 t=4.01...u are correct:)

OpenStudy (toxicsugar22):

A classroom has a volume of 288m3

OpenStudy (toxicsugar22):

What is its volume in km3

OpenStudy (shrutipande9):

i m not sure...ask in the physics group.

OpenStudy (toxicsugar22):

ok

OpenStudy (kayne):

Please post your questions in the appropriate group which is physics in this case.!